OpenStudy (anonymous):

Determine whether the vectors u and v are parallel, orthogonal, or neither. u = <6, -2>, v = <8, 24>

OpenStudy (jdoe0001):

get their dot product if their dot product is 0, they're orthogonal if one vector is just a multiple of the other... then they're parallel

OpenStudy (anonymous):

To get their dot product, do you multiply or add them together?

OpenStudy (jdoe0001):

actually... both you multiply them, and then you add them =) \(\bf <a,b>\cdot <c,d>\implies a\cdot c+b\cdot d\)

OpenStudy (anonymous):

so its <6,-2> x <8,24> (6 x 8) + (-2 x 24) <42,-48>

OpenStudy (jdoe0001):

well.... is supposed to give you a constant... no a vector \(\large \bf \bf <a,b>\cdot <c,d>\implies a\cdot c+b\cdot d\)

OpenStudy (anonymous):

so its just (6 x 8) + (-2 x 24) ? how does that work?

OpenStudy (jdoe0001):

yeap

OpenStudy (anonymous):

So how do I know if thay are parallel orthogonal or neither>

OpenStudy (jdoe0001):

get their dot product if their dot product is 0, they're orthogonal if one vector is just a multiple of the other... then they're parallel

OpenStudy (anonymous):

<42,-48> Is that the dot product?

OpenStudy (jdoe0001):

well.... is supposed to give you a constant... no a vector though \(\large \bf \bf <a,b>\cdot <c,d>\implies a\cdot c+b\cdot d\)

OpenStudy (anonymous):

So the vector is <42,-48>? or is it -6? which isnt zero or a multiple... So neither?

OpenStudy (jdoe0001):

hmmm recheck your dot product

OpenStudy (jdoe0001):

but the dot product gives a constant.. not a vector though

OpenStudy (anonymous):

<48, -48> equals zero orthogonal!

OpenStudy (jdoe0001):

yea 48 - 48 yes... is the dot product.. and that's 0, thus they're orthogonal, "perpendicular", then

OpenStudy (anonymous):

AH! Thank you! Quick question- when doing dot products with three numbers each, is it the same thing?

OpenStudy (jdoe0001):

yes

OpenStudy (anonymous):

Thank you!

OpenStudy (jdoe0001):

yw